### v3

# ConceptPsy: A Benchmark Suite with Conceptual Comprehensiveness in Psychology

Abstract
The critical field of psychology necessitates a comprehensive benchmark to enhance the evaluation and development of domain-specific Large Language Models (LLMs). Existing MMLU-type benchmarks, such as C-EVAL and CMMLU, include psychology-related subjects, but their limited number of questions and lack of systematic concept sampling strategies mean they cannot cover the concepts required in psychology. Consequently, despite their broad subject coverage, these benchmarks lack the necessary depth in the psychology domain, making them inadequate as psychology-specific evaluation suite. To address this issue, this paper presents  ConceptPsy, designed to evaluate Chinese complex reasoning and knowledge abilities in psychology.  ConceptPsyincludes 12 core subjects and 1383 manually collected concepts. Specifically, we prompt GPT-4 to generate questions for each concept using carefully designed diverse prompts and hire professional psychologists to review these questions. To help to understand the fine-grained performances and enhance the weaknesses, we annotate each question with a chapter label and provide chapter-wise accuracy. Based on  ConceptPsy, we evaluate a broad range of LLMs. We observe that, although some LLMs achieve similar accuracies on overall performances, they exhibit significant performance variations across different psychology concepts, even when they are models from the same series. We hope our work can facilitate the development of LLMs in the field of psychology.
11footnotetext: Equal Contribution.22footnotetext: Corresponding author.33footnotetext: Independent Researcher.
1 Introduction
Figure 1: GPT-3.5-Turbo’s concept-wise performance on Psychological Statistics. The x-axis represents the sequence of concepts, arranged in the order they appear in the textbook. The dashed circles represent sampled questions. It can be observed that when the tested concept in a sampled question only covers a subset, different samplings can mislead people’s understanding of a model. Although GPT-3.5-Turbo achieves an average score of 75% upon re-running, its performance across different concepts varies significantly.
Figure 2: Diagram overview of concepts in  ConceptPsy. We sample questions based on the requirement of the National Post-graduate Entrance Examination in China. Each question is tagged with a modified chapter name, serving as the chapter-level concept, to further provide chapter-level accuracy.
Domain benchmarks is crucial for the development and adaptation of Language Large Models (LLMs). Recently, there has been a significant advancement in LLMs, exemplified by models like GPT-4  (Achiam et al., 2023) , Claude-3 (Anthropic, 2023), Qwen2 (Bai et al., 2023a), DeepSeek-V2  (DeepSeek-AI, 2024). These models have shown extraordinary capabilities across a range of conventional tasks. However, their systematic understanding of key areas such as psychology remains unclear. This uncertainty hinders their effective adaptation to the psychology domain. Consequently, the effectiveness of evaluating these models through a well-designed benchmark becomes crucial for their further development.
While there are established English or Chinese evaluation benchmarks for foundational models such as MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2023), and C-EVAL (Huang et al., 2023), they present a limitation. Although these benchmarks encompass a broad range of subjects, they only include a few hundred questions per subject. This limitation is significant when adapting a model to a specific domain, as a thorough understanding within a domain is crucial for its development. The broad scope of these benchmarks doesn’t guarantee comprehensive coverage of necessary concepts, thereby limiting their effectiveness. Moreover, these benchmarks do not emphasize the comprehensive concept sampling when collecting questions. We discuss this in Section §2, highlighting the considerable concept bias in these benchmarks. This implies their questions might only cover a small subset of required concepts required in a domain, potentially leading to misleading results. For instance, as depicted in Figure 1, there’s a noticeable performance difference across different concepts for GPT-3.5-Turbo. While there are benchmarks designed for individual domains like Law  (Fei et al., 2023), Finance  (Zhang et al., 2023), Mathematics  (Wu et al., 2024), and Medicine  (Chen et al., 2024), a comprehensive evaluation benchmark for the field of Psychology, especially in the Chinese context—one of the world’s most widely spoken languages—is still absent.
To narrow the gap between the development of Chinese LLMs in psychology and their evaluation, we introduce  ConceptPsy, the first comprehensive Chinese evaluation suite designed to thoroughly assess LLMs’ psychology knowledge and reasoning abilities within a Chinese context. To ensure a comprehensive collection of concepts, we manually gather 1383 college-level concepts based on the official requirements of National Post-graduate Entrance Examination. Directly sourcing questions from textbooks or online question banks could raise copyright issues, potentially limiting the open-source nature of our benchmark. Consequently, based on the Chinese Academy of Sciences Psychological qualification exam, we carefully designed multiple types of prompts. Then we prompt the powerful GPT-4 to generate four multiple-choice questions for each concept. We hire the professional psychologists to review these questions, ensuring their correctness. Beyond overall accuracy, we also annotate a chapter-level label to each question to provide a fine-grained performance at the chapter level. These fine-grained results can assist developers in understanding the strengths and weaknesses of their models, thereby facilitating further improvements.
We carried out experiments to assess a broad range of LLMs using  ConceptPsy. Some strong Chinese models like Yi-34B-Chat and Qwen1.5-72B-Chat also achieve high overall scores. However, their performance disparities across different chapters can reach up to 50%. Even though some large-scale models achieve similar overall scores and belong to the same series, we observe their performance to vary significantly across different chapters, leading to analogous outcomes. We hope that  ConceptPsywill guide developers in understand the capabilities of their models in psychology from a comprehensive and fine-grained perspective and facilitate the growth of foundational psychology models for Chinese users.
Figure 3: Examples of concepts. We define a “concept” as fundamental units of understanding that encapsulate specific knowledge within a broader field of study.
Figure 4: Overview of Our Concept-Driven Framework. We collect relevant concepts based on the requirements of corresponding examinations. To diversify the types of questions, we summarize three question patterns from these exams and design specific prompts for each type. Questions are then generated using GPT-4. Subsequently, we hire professional psychological counselors to review the questions for accuracy and relevance.
Subject Coverage Rate
STEM
Computer Architecture 0.55
Computer Network 0.31
High School Biology 0.52
Social Science
High School Geography 0.48
Marxism 0.70
Humanity
High School History 0.76
Logic 0.47
Avg 0.54
Table 1: The concept coverage rate of randomly sampled subjects, with chapter-level concepts generated by GPT-4, as evaluated on C-EVAL.
Subject Benchmark #Questions #Concepts Coverage_rate
Professional Psychology C-EVAL - - -
CMMLU 232 84 (chapter-level) 0.59
Advanced Math C-EVAL 173 94 0.54
CMMLU 104 36 0.35
Table 2: The concept coverage rate of subjects with manually collected required concepts. We prompt GPT-4 to classify each question into one or more concepts and subsequently calculate the coverage rate.
2 Case Study on Chinese MMLU benchmarks
We conduct a case study to analysis the concept bias in popular Chinese MMLU benchmarks: C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023). We define a concept as a fundamental unit in the human learning process , and a chapter-level concept as the modified chapter title where a tested concept is located. For example, as illustrated in Figure 3 , the “random error” can serve as a concept when studying “random variables”. We study concept bias from two perspectives: 1. concept coverage rate (§2.1): the proportion of concepts tested in the sampled questions to the total concepts required in a subject. 2. performance variance (§2.2): the difference in model performance across various concepts.
2.1 Analysis on the Concept Coverage
We define the “concept coverage rate” as the ratio between the number of concepts $C_{collection}$ tested by the collected questions and the number of concepts $C_{requirement}$ required by the subject:$Concept\ coverage\ rate=\frac{C_{collection}}{C_{requirement}}$. In MMLU benchmarks, a question serves as an effective tool to assess a model’s knowledge comprehension and reasoning ability about the concepts being tested. Psychology is crucial for AI to comprehend human behavior. However, popular Chinese MMLU benchmarks, such as C-EVAL, do not include psychology. CMMLU includes a subject called “professional psychology” represented by 232 questions, and uses the average score to reflect a model’s ability in psychology. We study the concept coverage rate in psychology for CMMLU. We also sample various subjects from each discipline to explore the potential existence of concept bias in other subjcts.
Setup
Firstly, we calculate the concept coverage in psychology. We manually collect all college-level concepts (total 1383 concepts) based on the requirements of the National Post-graduate Entrance Examination in China. We deploy the robust GPT-4 to categorize each question into one or more concepts. Due to the constraints of context length, rather than categorizing each question into specific concepts, we utilize chapter-level concept tags for each question. It’s important to note that this approach may yield a higher score than the instance-level concept coverage rate. Given the scarcity of psychology-related subjects in the existing Chinese MMLU benchmark, we sampled some subjects from various disciplines (STEM, Social Science, Humanity) to further investigate the concept bias in current MMLU benchmarks. Manually collecting all required concepts for each subject is costly and impractical. Therefore, we randomly select advanced math and manually collect the concepts it requires. For the other sampled subjects, we carefully prompt GPT-4 to generate the first/second/third-level chapter names that represent the required concepts for each subject. We also prompted GPT-4 to categorize each question to chapters. We calculated the coverage rate of the lowest-level headings to represent the concept coverage rate. More experiments details can be found in Appendix D.
Subject Benchmark min max mean std
Professional Psychology C-EVAL - - - -
CMMLU 0.0 1.0 0.58 0.34
Advanced Math C-EVAL 0.0 0.80 0.35 0.24
CMMLU 0.0 0.75 0.41 0.18
Table 3: Performance of GPT-3.5-Turbo on subjects with manually collected required concepts across different chapter-level concepts.
Subject Baichuan2-13B Qwen1.5-MoE-A2.7B
Computer Architecture 0.50±0.17 0.72±0.14
Computer Network 0.59±0.19 0.80±0.17
High School Biology 0.52±0.25 0.68±0.32
High School Geography 0.73±0.16 0.84±0.19
Marxism 0.88±0.10 0.96±0.03
High School History 0.59±0.33 0.80±0.29
Logic 0.46±0.36 0.60±0.34
Table 4: The mean and standard deviation of model performance on randomly sampled C-EVAL subjects across different chapters.
The concept coverage rate is Low.
Although psychology is an vital domain for AI to learn human behavior, it is absent in C-EVAL, and its chapter-level (we use the chapter name instead of concept instance to calculate due to the budget limitation of GPT-4) concept coverage in CMMLU is only 59%. Questions with such an extremely low concept coverage rate could potentially mislead developers. Given the diverse range of subjects within psychology, it’s not surprising that the representation of a single subject in CMMLU is low. However, even for fields like advanced mathematics, which naturally focus on a single subject, the coverage rate remains low.
Compared to CMMLU, which uses general task names (e.g., computer science), C-EVAL employs more specific subject names, which facilitates the collection of accurate required concepts. Therefore, we further calculate the coverage rate of different subjects within various disciplines in C-EVAL (Table 1). High School History and Marxism have relatively high coverage rates, but they only reach around 70%. The average coverage rate is only 54%, which might be lower than what users expect based on the average scores reflecting the number of concepts in these subjects.
2.2 Variations in Performance Across Different Concepts
Setup
Based on the questions categorized into various chapter-level concepts as discussed in §2.1, we further calculate the variance in model performance across these different concepts.
The various of performance across different chapters is large.
A high variance across different concepts can exacerbate the impact of a low concept rate on the average score, making it more misleading. As illustrated in Table 3 and 4, strong models like GPT-3.5-Turbo(OpenAI, 2022) exhibit a standard deviation exceeding 10% across different chapters. We also test open-source models such as Baichuan2-13B(Yang et al., 2023a) and Qwen1.5-MoE-A2.7B (Bai et al., 2023b). They continue to demonstrate a standard deviation of over 10% on subjects sampled from different disciplines.
3  ConceptPsy
Category # C # Q $\mathbf{L_{Q}}$ $\mathbf{L_{A}}$
In terms of subject
Clinical & Counseling Psychology 56 156 38.9 12.0
Psychology of Personality 91 318 36.7 11.1
Abnormal Psychology 89 268 35.8 12.4
History of Psychology 126 472 27.4 10.5
General Psychology 183 605 35.3 9.4
Psychometrics 115 368 43.2 10.1
Social Psychology 169 559 26.2 13.7
Management Psychology 88 315 37.2 10.4
Psychological Statistics 99 311 57.0 8.4
Experimental Psychology 141 413 59.3 9.3
Developmental Psychology 159 580 30.7 11.3
Educational Psychology 67 208 42.9 11.5
In terms of split
Dev - 60 - -
Valid - 428 - -
Test - 4085 - -
Total 1383 4573 - -
Table 5: Statistics of  ConceptPsy. The column “#C” indicates the number of concepts we have annotated for each subject, with each concepts generating 4 questions and filtered by professional psychological annotators. The number of questions obtained after the review process is displayed in the column “#Q”. $L_{Q}$ and $L_{A}$ is the average length of a question and answer separately.
For popular Chinese MMLU benchmarks, C-EVAL lacks the critical field in human society: psychology. While CMMLU reflects the model’s related ability in psychology with an average score derived from 232 questions, its chapter-level (84 chapters) concept coverage rate is only 59%, the instance-level (1383 instances ). These findings highlight the urgent need for an MMLU-style psychology benchmark with low concept bias. Instead of only providing an average accuracy, we also label each question a chapter-level concept to provide more fine-grained results.
3.1 Overview
[…]
Figure 5: The review rules the question review process are as follows. Professional psychological annotators will filter, modify, and review the generated questions based on these requirements.
  • Low Concept Bias:  ConceptPsycovers all required concepts according to the graduate school entrance examination for the 12 core subjects in psychology.
  • Fine-grained Score: Instead of only providing an average score for the whole subject, we also provide a chapter-level concept score.
  • Focus on Psychology:  ConceptPsy  concentrates on the psychology domain, aiming to tackle the absence and concept bias issues in psychology within previous MMLU benchmarks.
3.2 Data Collection
Subject Selection
Our selection of 12 core standard subjects (shown in Table 5) within the discipline of psychology has been meticulously considered in accordance with both higher education standards and professional qualification requirements. We select 11 courses from the Peking University Core Curriculum Handbook for Undergraduate Program. Additionally, we included Psychological Counselling as a fundamental subject within our data set to adhere to professional qualification criteria.
Concepts Collection
The Graduate Entrance Examination is an official test that evaluates the professional knowledge acquired by undergraduate students. We collect concepts based on the requirement of the 12 subjects in this exam. An example of our method for recording concept is depicted in Figure 6. To improve the quality of generated questions, we provide the GPT-4 model not only with the concepts but also with the first/second level chapter names.
To generate high-quality questions, we meticulously design prompts by analyzing professional exam question types and categorizing them into three groups: (1) Calculation; (2) Theory understanding; (3) Case Study. As shown in Figure 6, we assign an appropriate question type based on the category of the knowledge point during question generation. We utilize four types of prompts (details can be found in Appendix H). Three of these prompts are used to generate the three types of questions, and the fourth type is used to generate all three types of questions for the same concept simultaneously. Figure 14 provides an example of a calculation type multiple-choice question. To ensure a relatively even number of questions for each concept, we generate four questions per concept.
Figure 6: An example of an annotator assigning a suitable prompt to a concept. For the concept “random error”, we collect multiple descriptions. The appropriate prompt is assigned based on the type of description provided.
Clinical &
Counseling
Psy of
Personality
Abnormal
Psy
History of
Psy
General
Psy
Psy-
chometrics
Social
Psy
Management
Psy
Psychological
Statistics
Experimental
Psy
Developmental
Psy
Educational
Psy
Avg
Chinese-Alpaca-2-7B 0.59 0.58 0.59 0.52 0.45 0.51 0.63 0.62 0.46 0.58 0.59 0.67 0.57
Llama-2-13B-Chat 0.66 0.62 0.65 0.56 0.54 0.57 0.66 0.63 0.48 0.57 0.58 0.66 0.6
Chatglm2-6B 0.71 0.63 0.68 0.63 0.62 0.62 0.69 0.64 0.54 0.64 0.66 0.72 0.65
Llama-2-70B-Chat 0.74 0.66 0.73 0.62 0.61 0.59 0.75 0.7 0.52 0.63 0.69 0.69 0.66
Mistral-7B-Instruct 0.73 0.65 0.66 0.59 0.62 0.63 0.71 0.72 0.52 0.65 0.66 0.74 0.66
Baichuan2-7B-Chat 0.72 0.69 0.77 0.65 0.67 0.66 0.75 0.72 0.53 0.66 0.7 0.76 0.69
Baichuan2-13B-Chat 0.78 0.73 0.8 0.68 0.73 0.69 0.82 0.75 0.6 0.71 0.74 0.83 0.74
GPT-3.5-Turbo 0.78 0.74 0.8 0.67 0.73 0.7 0.82 0.8 0.64 0.74 0.77 0.81 0.75
Mixtral-8x7B-Instruct 0.82 0.76 0.78 0.65 0.73 0.7 0.83 0.78 0.71 0.73 0.74 0.79 0.75
Qwen1.5-7B-Chat 0.83 0.75 0.84 0.66 0.75 0.67 0.79 0.81 0.61 0.72 0.73 0.83 0.75
Internlm2-7B-Chat 0.86 0.75 0.84 0.68 0.78 0.7 0.82 0.82 0.62 0.75 0.78 0.81 0.77
Yi-6B-Chat 0.85 0.78 0.89 0.71 0.84 0.7 0.85 0.86 0.62 0.72 0.78 0.83 0.79
Qwen1.5-72B-Chat 0.87 0.74 0.89 0.8 0.81 0.77 0.88 0.87 0.38 0.77 0.83 0.85 0.79
Qwen1.5-14B-Chat 0.86 0.8 0.85 0.72 0.83 0.72 0.83 0.85 0.67 0.81 0.79 0.85 0.8
Yi-34B-Chat 0.92 0.84 0.91 0.81 0.86 0.78 0.88 0.9 0.69 0.84 0.86 0.86 0.85
Table 6: Performances on  ConceptPsywith different LLMs.
3.3 Questions Review
After using GPT-4 to generate the questions, three professional psychological annotators will review each one to ensure the quality of the dataset. We primarily review the correctness of the questions and filter out those that are unreasonable. The rules and precautions for question review are shown in Figure 5. The annotators will check the rationality of the knowledge points, the questions, and their relationship. If there are errors in a question, modifications require a consensus from at least three psychological counselors to reduce subjective bias and mistakes.
4 Experiments
Figure 7: Concept-level results for models more than 34B, in addition to GPT-3.5-Turbo.
4.1 Dataset Statistics
We carefully select 12 psychology subjects, including: clinical and counseling psychology, psychology of personality, abnormal psychology, history of psychology, general psychology, psychometrics, social psychology, management psychology, psychological statistics, experimental psychology, developmental psychology, educational psychology. We manually collected 1383 concept instances from the National Post-graduate Entrance Examinationdataset. We generated 4 questions per instance, resulting in 4573 high-quality questions after a rigorous review and filtering process. Additionally, we collected 84 chapter-level concepts and assigned labels to each question to evaluate chapter-level performance.
4.2 Setup
We evaluate a diverse set of strong models to ensure a comprehensive evaluation. More details about the models can be found in Appendix G . We conduct the evaluations using a 5-shot method. We set the temperature as 0, $top\_p$ as 1.0.
4.3 Results
In Table 6, we showcase the performance of popular Chinese LLMs. The Yi-34B-Chat model outperforms, achieving an average score of 0.85, surpassing strong models such as Qwen1.5-72B-Chat and GPT-3.5-Turbo. Both Qwen1.5-14B-Chat and Qwen1.5-72B-Chat exhibit similar performance, which may suggest a lack of training data in the psychology domain, preventing the models from leveraging their larger parameter counts. Many smaller Chinese models, such as Yi-6B-Chat and Qwen1.5-7B-Chat, outperform larger and stronger models like Mixtral-8x7B-Instruct, potentially due to the former having more training data in Chinese. Most models demonstrate lower performance in Psychological Statistics and Psychometrics compared to other subjects, implying these models need to enhance their reasoning abilities in psychology.
4.4 Chapter-level Performance
ConceptPsyalso provides chapter-level performance insights, offering more detailed information. From Figure 7, we can understand why the strong Qwen-72B-Chat model’s average score is lower than the smaller model Yi-34b-Chat. Qwen-72B-Chat underperforms Yi-34b-Chat in most chapters, particularly in reasoning chapters such as Applied Calculations in Statistical Psychology or Statistical Charts. In contrast, Mistral-8x7B-Instruct performs worse in theoretical chapters, possibly indicating a lack of corresponding knowledge in its training data. All models perform poorly in non-parametric tests, suggesting potential areas of improvement for developers. More chapter-level results can be found in Appendix E
5 Related Works
5.1 Chinese MMLU benchmarks
While the evolution of English language benchmarks continues to flourish (Hendrycks et al., 2020; Huang et al., 2023; Li et al., 2023; Zhong et al., 2023), Chinese MMLU benchmarks remain underdeveloped. The CLUE benchmark (Xu et al., 2020) is a widely used large-scale NLU benchmark for Chinese.The AGIEval benchmark (Zhong et al., 2023) expands this with questions from various Chinese exams, while MMCU (Zeng, 2023) includes questions from diverse domains. MCU (Zeng, 2023) expands on this by collecting questions from a broader range of domains. The C-EVAL benchmark Huang et al. (2023) gathers questions from different levels, including middle school, high school, and professional qualification exams, employing strategies to mitigate dataset leakage by using non-paper-based questions and simulation questions. ConceptMath (Wu et al., 2024) also evaluates LLM’s ability in a concept-wise manner. However, it differs from our work in that it (1) focuses on mathematics, (2) is not a comprehensive MMLU benchmark, and (3) is released more than three months after our work.
5.2 Benchmarks for LLMs in psychology
Although psychology is a crucial domain for achieving artificial intelligence, there are currently few benchmarks in this field. CMMLU has collected hundreds of questions as "professional psychology" to evaluate a model’s knowledge understanding and reasoning capabilities. Other works (Yang et al., 2023b; Amin et al., 2023; Lamichhane, 2023) approach this from a mental health perspective, utilizing various classification tasks, including various emotions, suicidal tendencies, and more. PsyEval (Jin et al., 2023) further expands in the mental health field by adding tasks such as Diagnosis Prediction.
6 Discussion
This paper introduces  ConceptPsy, the first comprehensive benchmark for manually collected psychology concepts in the psychology domain. This addresses the current issue where popular Chinese MMLU benchmarks either lack psychology subjects or have significant concept bias in psychology. In addition to providing an average score for each subject in  ConceptPsy, we manually label each question with a chapter-level concept to offer more fine-grained results. This assists developers in better improving and deploying their models. Furthermore, we conduct a case study on the presence of concept bias in current Chinese MMLUs, further underscoring the urgent need for  ConceptPsy.
Limitations
We have solely explored the concept-related bias within the Chinese MMLU series of datasets and have not investigated biases in other areas. Additionally, due to resource constraints, we have only collected and analyzed the concepts from limited subjects. Our work exclusively examines Chinese benchmarks and does not extend to benchmarks in other languages.
Ethics Statement
We have thoroughly examined our data to ensure that there are no ethical issues. The data is generated by GPT-4, using concepts prescribed by the National Entrance Examination for Postgraduates. Furthermore, the generated data is subjected to rigorous scrutiny by professional psychologists to ensure its ethical soundness.
Appendix A Human Evaluation
Subjects Pro Non-Pro
Clinical & Counseling Psychology 0.85 0.28
Psychology of Personality 0.82 0.24
Abnormal Psychology 0.89 0.32
History of Psychology 0.73 0.22
General Psychology 0.9 0.26
Psychometrics 0.75 0.3
Social Psychology 0.82 0.16
Management Psychology 0.8 0.36
Psychological Statistic 0.78 0.2
Experimental Psychology 0.82 0.32
Developmental Psychology 0.83 0.16
Educational Psychology 0.82 0.16
Avg 0.82 0.25
Table 7: Human evaluation results of  ConceptPsy. “Pro” and “Non-pro” represent Professional Counselor and Non-Professional Student respectively.
We also provide human performance on the subset of  ConceptPsy. Specifically, we randomly sample 50 questions from each of 12 subjects, totaling 600 questions. A professional counselor independently completes all 600 questions, while ten non-experts are each assigned different subjects for evaluation. The results are shown in Table 7. The professional counselor significantly outperforms GPT-3.5, achieving an average score of 82%. However, the performance of non-expert undergraduates and graduate students is poor, averaging 25%, which is close to random. This underscores our benchmark’s demand for specialized knowledge in psychology and its effectiveness in distinguishing whether a model has mastered these professional concepts.
Subjects Win Rate Tie Rate
Clinical & Counseling Psychology 0 0.8
Psychology of Personality 0 0.9
Abnormal Psychology 0 0.8
History of Psychology 0.1 0.8
General Psychology 0.1 0.9
Psychometrics 0.1 0.7
Social Psychology 0.82 0.16
Management Psychology 0.1 0.8
Psychological Statistic 0.4 0.5
Experimental Psychology 0.2 0.7
Developmental Psychology 0.4 0.4
Educational Psychology 0.3 0.7
Avg 0.27 0.62
Table 8: We hire two professional counselor to compete the quality of generated questions and human-designed questions.
Appendix B Competition with Human-designed Questions
We randomly selected 15-20 questions for each subject and an equivalent number of questions from real exams available online. We shuffle the order of the questions, presenting one generated and one real question side by side, and ask two psychological annotators to judge which was better or if there was a tie. The averaged results are shown in Table 8. The “Win Rate” indicates the proportion of instances where our generated questions were rated better than the real ones. As can be seen, the majority of the generated questions matched or even surpassed the performance of actual exam questions. This success can be attributed to two factors: 1) our prompts are a carefully designed summary of the question types found in Chinese psychological professional examinations; 2) GPT-4’s robust capabilities in synthetic data.
Appendix C Chapter-level Statistic of ConceptPsy
In Table 9 and Table 10, we present the chapters of each subject along with the corresponding number of concepts and questions.
Subject Chapter # C # Q
Clinical & Counseling Psychology History of Clinical and Counseling Psychology 4 13
Basic concepts of psychotherapy and counseling 3 7
Characteristics of the therapeutic relationship and its influences 10 32
Work Ethics in Clinical and Counseling Psychology 5 16
Research Methods in Clinical and Counseling Psychology 8 25
Theory and Practice of Counseling and Psychotherapy 25 63
Psychology of Personality Basic Concepts of Personality Psychology 12 43
Psychoanalytic School 23 84
Behaviorist school of personality theory 10 37
Cognitive School 9 33
Humanistic School 11 40
Personality Trait Theory 18 57
Biological School 4 13
Positive Psychology 4 11
Abnormal Psychology Basic Concepts of Abnormal Psychology 17 60
Anxiety Disorders 26 81
Mood Disorders 7 19
Eating Disorders 6 18
Personality Disorders 14 40
Substance Dependence 10 28
Childhood Mental Disorders 8 22
History of Psychology The Origins and Establishment of Western Psychology 17 62
Psychology of Consciousness 40 154
Behaviourist Psychology 21 82
Psychoanalysis 25 93
Cognitive Psychology 13 48
Humanistic Psychology 8 30
History of Chinese Psychology 1 3
General Psychology Basic Concepts in General Psychology 15 48
Biological Basis of Mind and Behaviour 9 33
Consciousness and Attention 15 51
Sensation 15 46
Perception 17 60
Memory 16 53
Thinking 21 67
Speech 7 21
Moods and Emotions 15 50
Motivation, Needs & Will 15 52
Competencies 15 50
Personality Theory 22 74
Psychometrics Basic Concepts in Measurement Psychology 12 41
Classical Measurement Theory 24 91
Basic Concepts of Psychological Testing 40 105
Commonly Used Psychological Tests 39 129
Table 9: Details of the number of Concepts and Questions in each Chapter.
Subject Chapter # C # Q
Social Psychology History of Social Psychology 40 134
Social Thinking 48 163
Social Relations 27 94
Social Influence 24 79
Basic Concepts of Social Psychology 30 89
Management Psychology Management Philosophy 18 65
Organizational Motivation 18 80
Leadership Theory 18 65
Organizational Theory 33 105
Psychological Statistics Basic Concepts in Statistical Psychology 3 12
Statistical Charts 5 15
Concentration 3 13
Measures of Variation 7 34
Correlation 2 5
Mathematical Basis of Inferential Statistics 16 65
Parameter Estimation 13 28
Hypothesis Testing 13 53
Applied Computing in Statistical Psychology 21 62
Chi-Square Tests 2 5
Non-Parametric Tests 2 8
Preliminary Multivariate Statistical Analysis 2 11
Experimental Psychology Basic Concepts of Experimental Psychology 3 12
Variables in Psychological Experiments 23 75
Design of Psychological Experiments 12 38
Reaction Time Method 9 33
Psychophysical Methods 40 97
Major Psychological Experiments 47 158
Developmental Psychology Basic Concepts in Developmental Psychology 5 24
Basic Theories of Psychological Development 19 90
Biological Basis of Psychological Development and Fetal Development 9 33
Intelligence 9 34
Emotions 8 34
Early Childhood Psychological Development 36 114
Child Psychological Development 27 90
Adolescent Psychological Development 18 60
Psychological Development in Adulthood 27 101
Educational Psychology Basic Concepts of Educational Psychology 6 18
General Psychology of Learning 6 21
Major Theories of Learning 29 100
Categorical Learning Psychology 25 69
Table 10: Details of the number of Concepts and Questions in each Chapter (Cont.)
Appendix D Experiments Details of Calculating Concept Coverage Rate
To calculate the concept coverage rate, the challenge lies in obtaining the required concepts in a subject. For advanced math, we first collect its chapters and the concepts under different chapters based on relevant exam requirements. We initially prompt GPT-4 to classify each question into different chapters, then prompt GPT-4 to further classify the question into specific instance-level concepts. We calculate the concept coverage rate and performance variance based on the concepts covered by these questions. This hierarchical classification allows for more accurate categorization.
For psychology in CMMLU, due to the vast number of concepts, prompting GPT-4 with these concepts in the input prompt each time is too expensive, so we use chapter-level concepts. We collect 84 chapter names according to the requirements, and classify each question into one or more chapters.
For other subjects, manually collecting concepts is too costly, and it’s impossible to collect concepts for each subject. We prompt GPT-4 to generate a 3-level syllabus for the subject as the required concepts. We then prompt GPT-4 to categorize each question into one or more first-level headings in the syllabi. We further classify them into the second- and third-level headings under the selected first-level headings.
Appendix E More Results on Fine-grained Performance
We provide more chapter-level results in this Figure 8, 9.
Figure 8: Concept-level results for models wi 14 Billion parameters
Figure 9: Concept-level results for some other models
Appendix F Qualifications of Our psychology Annotators
Our psychological annotators is professional. We enlist three psychological annotators with diverse expertise. The first is a postdoctoral researcher specializing in experimental and counseling psychology. The second is a registered psychology researcher with the British Psychological Society with five years of research experience in interdisciplinary psychology laboratories. The third, a counseling psychologist, brings over three years of practical counseling experience.
Appendix G Details About Evaluated Models
Model Name Model Code/API
Chinese-Alpaca-2-7BCui et al. (2023) hfl/chinese-alpaca-2-7b
Chatglm2-6BDu et al. (2022) THUDM/chatglm2-6b
Llama-2-13B-ChatTouvron et al. (2023) meta-llama/Llama-2-13b-chat-hf
Llama-2-70B-ChatTouvron et al. (2023) meta-llama/Llama-2-70b-chat-hf
Baichuan2-7B-ChatYang et al. (2023a) baichuan-inc/Baichuan2-7B-Chat
Baichuan2-13B-ChatYang et al. (2023a) baichuan-inc/Baichuan2-13B-Chat
Mistral-7B-Instruct-v0.2Jiang et al. (2023) mistralai/Mistral-7B-Instruct-v0.2
Mixtral-8x7B-Instruct-v0.1Jiang et al. (2024) mistralai/Mixtral-8x7B-Instruct-v0.1
Qwen1.5-7B-ChatBai et al. (2023b) Qwen/Qwen1.5-7B-Chat
Qwen1.5-72B-ChatBai et al. (2023b) Qwen/Qwen1.5-72B-Chat
Qwen1.5-14B-ChatBai et al. (2023b) Qwen/Qwen1.5-14B-Chat
Internlm2-7B-ChatCai et al. (2024) internlm/internlm2-chat-7b
Yi-6B-ChatAI et al. (2024) 01-ai/Yi-6B-Chat
Yi-34B-ChatAI et al. (2024) 01-ai/Yi-34B-Chat
GPT-3.5-TurboOpenAI (2022) Azure api: gpt-35-turbo
Table 11: Model code/API of our evaluated models.
The evaluated models include Chinese-Alpaca-2-7BCui et al. (2023), Chatglm-6BDu et al. (2022), Chatglm2-6BDu et al. (2022), Llama-2-13B-ChatTouvron et al. (2023), Llama-2-70B-ChatTouvron et al. (2023), Baichuan2-7B-ChatYang et al. (2023a), Baichuan2-13B-ChatYang et al. (2023a), Mistral-7B-Instruct-v0.2Jiang et al. (2023), Mixtral-8x7B-Instruct-v0.1Jiang et al. (2024), Qwen1.5-7B-ChatBai et al. (2023b), Qwen1.5-72B-ChatBai et al. (2023b), Qwen1.5-14B-ChatBai et al. (2023b), Internlm2-7B-ChatCai et al. (2024), Yi-6B-ChatAI et al. (2024), and Yi-34B-ChatAI et al. (2024). The model codes can be found in Table 11 . We evaluate these models with vLLM (Kwon et al., 2023).
Appendix H Prompts for Questions Generation
We totally design four distinct prompts to steer GPT-4 in generating questions based on provided knowledge points. In designing these prompts, we have taken into account several guidelines. Firstly, the generated questions should exhibit a high level of difficulty and complexity. Secondly, while generating questions, GPT-4 should primarily rely on the given knowledge points, but it can also incorporate its inherent psychological knowledge. Thirdly, each knowledge point unit may yield multiple questions, but the content, type, or perspective of the questions should be distinct.
Figures 10, 11, 12, and 13 show the specific prompts that we have inputted into GPT-4. These prompts have been meticulously designed, with each one being tailored to control the generation of a different kind of question. The first three prompts correspond to generating Theory understanding, Case study, and Calculation type questions, respectively, while the last one encompasses all of the aforementioned types. We choose the appropriate prompt for each knowledge point based on its content.
Figure 10: Theory understanding
Figure 11: Case Study
Figure 12: Calculation
Figure 13: multiple type prompt
Figure 14: The question generation prompt template (translated in English), which is primarily designed for generating the type of calculation questions.
Figure 15: An example of prompts in few-shot setting. The black text is what we feed into model, while the red text is the response completed by model. The English translation for the Chinese input is provided in the purple text, which is not included in the actual prompt.