### v4

# ConceptPsy: A Benchmark Suite with Conceptual Comprehensiveness in Psychology

Abstract
The effective incorporation of Large Language Models (LLMs) into the field of psychology necessitates a comprehensive domain-specific benchmark to guide their development and adaptation. Existing Chinese benchmarks in the style of MMLU, such as C-EVAL and CMMLU, do include psychology subjects, but their concept coverage is far from exhaustive. The number of questions in each domain is just in the hundreds, and an uneven question sampling process can lead to a “concept bias” issue. This bias, stemming from using a question set with a low concept coverage rate to represent a subject, can potentially lead to skewed results. To address this, we present  ConceptPsy, a unique benchmark specifically designed for evaluating Chinese LLMs’ complex reasoning and knowledge in psychology.  ConceptPsy  encompasses 12 core subjects and 1,383 concepts from official exams. To avoid copyright issues, we prompt GPT-4 to generate questions for each of the concepts, which are then validated by psychology professionals to ensure high quality. Besides the overall scores, we annotate each question with a chapter label to provide fine-grained results. We evaluate a range of LLMs on  ConceptPsy  and the results show significant performance differences across psychology concepts, even among models from the same series. We anticipate the comprehensive concept coverage and the fine-grained strengths and weaknesses identified by  ConceptPsy  can facilitate the development and growth of the Chinese psychology domain.
11footnotetext: Equal Contribution.22footnotetext: Corresponding author.33footnotetext: Independent Researcher.
1 Introduction
Figure 1: GPT-3.5-Turbo’s concept-wise performance on Psychological Statistics. The x-axis represents the sequence of concepts, arranged in the order they appear in the textbook. The dashed circles represent sampled question sets. Different samplings can mislead people’s understanding of a model.
Figure 2: Diagram overview of concepts in  ConceptPsy. We sample questions based on the requirement of the National Post-graduate Entrance Examination in China. Each question is tagged with a modified chapter name, serving as the chapter-level concept, to further provide chapter-level accuracy.
Domain-specific benchmarks are essential for the advancement and fine-tuning of LLMs. Recently, there has been a significant advancement in LLMs, exemplified by models like GPT-4  (Achiam et al., 2023), Claude-3 (Anthropic, 2023), Qwen2 (Bai et al., 2023a), DeepSeek-V2  (DeepSeek-AI, 2024). These models have shown remarkable abilities in a wide array of standard tasks. However, a critical gap remains in the systematic understanding of these models’ performance within the important domain of psychology. This gap primarily arises from the absence of a comprehensive psychological benchmark.
While some Chinese MMLU-style benchmarks such as CMMLU (Li et al., 2023) have offered a wide range of subjects for assessment. They tend to focus more on covering many subjects rather than diving deep into each one. This is reflected by the limited number of questions per subject and the lack of emphasis on concept coverage during the sampling process. As a result, these questions do not thoroughly cover the necessary concepts within each subject. We term this problem as “concept bias” and discuss more in §2.1. Relying on question sets with limited conceptual breadth to represent a subject can lead to misleading results. Figure 1 illustrates the performance variability of models across concepts, with each question set distinguished by a unique color. These variations highlight that models can exhibit a range of proficiency—from poor to fair to proficient—based on the specific questions used for evaluation. As a result, it might mislead developers about how well a model can handle psychology-related topics.
Domain benchmarks are key to advancing LLMs and have been introduced in fields like (Fei et al., 2023), Finance (Zhang et al., 2023), Mathematics (Wu et al., 2024), Medicine (Chen et al., 2024). However, a comprehensive Chinese benchmark for evaluating advanced knowledge and reasoning abilities in psychology is missing. To narrow the gap, we introduce  ConceptPsy, the first comprehensive Chinese evaluation suite designed to thoroughly assess LLMs’ psychology knowledge and reasoning abilities. To ensure the comprehensiveness of concepts, we manually gathered 1383 college-level concepts based on the official requirements of National Post-graduate Entrance Examination. To avoid copyright issues, we design prompts inspired by the Chinese Academy of Sciences Psychological qualification exam. We prompt GPT-4 to generate four multiple-choice questions per concept, which are reviewed by professional psychologists. We also assigned a chapter-level label to each question, providing fine-grained performance analysis to facilitate model improvements.
We conducted experiments evaluating a wide array of LLMs. Results show GPT-4 surpassing human performance with an average accuracy of 84%. Chinese models like Yi-34B-Chat and Qwen1.5-72B-Chat also scored highly, though disparities in performance across chapters reached up to 50%. Even within the same series, large-scale models displayed significant chapter-specific performance variations. We believe  ConceptPsy  will help developers better comprehend their models’ psychology abilities and foster the development of foundational psychology models for Chinese users.
Figure 3: Examples of concepts. We define a “concept” as fundamental units of understanding that encapsulate specific knowledge within a broader field of study.
Figure 4: Overview of Our Concept-Driven Framework. We collect relevant concepts based on the requirements of corresponding examinations. To diversify the types of questions, we summarize three question patterns from these exams and design specific prompts for each type. Questions are then generated using GPT-4. Subsequently, we hire professional psychological counselors to review the questions for accuracy and relevance.
Subject Coverage Rate
STEM
Computer Architecture 0.55
Computer Network 0.31
High School Biology 0.52
Social Science
High School Geography 0.48
Marxism 0.70
Humanity
High School History 0.76
Logic 0.47
Avg 0.54
Table 1: The concept coverage rate of randomly sampled subjects, with chapter-level concepts generated by GPT-4, as evaluated on C-EVAL.
Subject Benchmark #Questions #Concepts Coverage_rate
Professional Psychology C-EVAL - - -
CMMLU 232 84 (chapter-level) 0.59
Advanced Math C-EVAL 173 94 0.54
CMMLU 104 36 0.35
Table 2: The concept coverage rate of subjects with manually collected required concepts. We prompt GPT-4 to classify each question into one or more concepts and subsequently calculate the coverage rate.
2 Case Study of Concept Bias
We conduct a case study to analyze the concept bias in popular Chinese MMLU benchmarks C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023). We define a concept as a fundamental unit in the human learning process. For example, in Figure 3, the “random error” is a concept when studying “random variables”. We study concept bias from two perspectives: 1. concept coverage rate (§2.1): the proportion of concepts tested in the sampled questions to the total concepts required. 2. performance variance (§2.2): the difference in model performance across various concepts.
2.1 Analysis on the Concept Coverage
In our research, we introduce the term “concept coverage rate” to quantify the extent to which the tested concepts in a given question set align with the required concepts in a given subject. Specifically, we calculate it as the ratio of the number of tested concepts, denoted as $C_{\text{collection}}$, to the number of required concepts, denoted as $C_{\text{requirement}}$: $(Concept_{coverage\_rate}=\frac{C_{collection}}{C_{requirement}}$. C-EVAL excludes subjects related to psychology. Within CMMLU, a subject termed “professional psychology" contains 232 questions, and the average score is utilized to gauge a model’s proficiency in psychology. Our analysis delves into the concept coverage rate specific to psychology within CMMLU. Additionally, we conduct sampling across diverse disciplines to investigate the potential presence of concept bias in other subjects.
Setup
We first assess the concept coverage of “professional psychology” in C-MMLU. Utilizing National Post-graduate Entrance Examination, we manually gather all necessary college-level concepts (1383 in total). We then employ GPT-4 to assign each question to one or more concepts. To investigate concept bias in other subjects, we sample subjects from various disciplines (STEM, Social Science, Humanities) in C-EVAL. Due to the resource limitation, we only manually collect the required concepts for advanced math. For other subjects, we use GPT-4 to generate chapter names representing the required concepts. Each question is also categorized into chapters using GPT-4. More experimental details can be found in Appendix E.
Subject Benchmark min max mean std
Professional Psychology C-EVAL - - - -
CMMLU 0.0 1.0 0.58 0.34
Advanced Math C-EVAL 0.0 0.80 0.35 0.24
CMMLU 0.0 0.75 0.41 0.18
Table 3: Performance of GPT-3.5-Turbo on subjects with manually collected required concepts across different chapter-level concepts.
Subject Baichuan2-13B Qwen1.5-MoE-A2.7B
Computer Architecture 0.50±0.17 0.72±0.14
Computer Network 0.59±0.19 0.80±0.17
High School Biology 0.52±0.25 0.68±0.32
High School Geography 0.73±0.16 0.84±0.19
Marxism 0.88±0.10 0.96±0.03
High School History 0.59±0.33 0.80±0.29
Logic 0.46±0.36 0.60±0.34
Table 4: The mean and standard deviation of model performance on randomly sampled C-EVAL subjects across different chapters.
The concept coverage rate is Low.
Although psychology is a vital domain for AI to learn human behavior, it is absent in C-EVAL, and its chapter-level concept coverage rate in CMMLU is only 59%. Questions with such an extremely low concept coverage rate could potentially mislead developers. This low coverage rate, although partly attributable to the broad spectrum of subjects within psychology, is not exclusive to this field. Even in more focused domains like advanced mathematics, the coverage rate is disappointingly low. Other subjects in C-EVAL (as shown in Table 1), such as High School History and Marxism, fare slightly better with coverage rates around 70%. However, the average coverage rate across all subjects is only 54%. This figure might fall short of user expectations, considering these questions are used as a representation of a subject.
2.2 Variations in Performance Across Different Concepts
A high variance in scores across concepts can make the average accuracy more misleading. To explore this, we measure the variance of different concepts within a subject.
Setup
Following the method in section §2.1, we categorize questions into chapter-level concepts and calculate the chapter-level performance variance.
The performance across different chapters varies greatly.
As demonstrated in Table 3 and 4, powerful models like GPT-3.5-Turbo (OpenAI, 2022) exhibit a standard deviation exceeding 10% across different chapters. We also evaluate open-source models such as Baichuan2-13B (Yang et al., 2023a) and Qwen1.5-MoE-A2.7B (Bai et al., 2023b). These models persist in showing a standard deviation of over 10% on subjects sampled from diverse disciplines. This high variance indicates that the final accuracy of these models is greatly influenced by the question set sampled.
3  ConceptPsy
Category # C # Q $\mathbf{L_{Q}}$ $\mathbf{L_{A}}$
In terms of subject
Clinical & Counseling Psychology 56 156 38.9 12.0
Psychology of Personality 91 318 36.7 11.1
Abnormal Psychology 89 268 35.8 12.4
History of Psychology 126 472 27.4 10.5
General Psychology 183 605 35.3 9.4
Psychometrics 115 368 43.2 10.1
Social Psychology 169 559 26.2 13.7
Management Psychology 88 315 37.2 10.4
Psychological Statistics 99 311 57.0 8.4
Experimental Psychology 141 413 59.3 9.3
Developmental Psychology 159 580 30.7 11.3
Educational Psychology 67 208 42.9 11.5
In terms of split
Dev - 60 - -
Valid - 428 - -
Test - 4085 - -
Total 1383 4573 - -
Table 5: Statistics of  ConceptPsy. The column “#C” indicates the number of concepts we have annotated for each subject, with each concepts generating 4 questions and filtered by professional psychological annotators. The number of questions obtained after the review process is displayed in the column “#Q”. $L_{Q}$ and $L_{A}$ is the average length of a question and answer separately.
Our proposed  ConceptPsy  seeks to fill the gap by offering a complete benchmark for evaluating the knowledge and reasoning ability of models in psychology. To the best of our knowledge, we are the first to explain the issue of concept bias in Chinese MMLU-style benchmarks and mitigate this problem by constructing question sets that cover necessary concepts.
3.1 Overview
In  ConceptPsy, we manually summarize 1383 concepts from the National Post-graduate Entrance Examination, then prompting GPT-4 to generate questions for each to ensure comprehensive coverage of all key concepts. This method sets  ConceptPsy  apart from prior Chinese MMLU benchmarks like CMMLU and C-EVAL, which relied on questions from online sources or textbooks. Our choice to use GPT-4 for question generation was driven by two main reasons: 1) To avoid copyright issues associated with using existing questions, facilitating broader industrial application; 2) Many concepts lacked associated multiple-choice questions, posing a challenge for creating new ones. To maintain question quality, we develop specific prompts based on the Professional Counselor Examination formats and had them reviewed by three professional psychologists.  ConceptPsy  differs from CMMLU and C-EVAL in the following ways:
Figure 5: An example of an annotator assigning a suitable prompt to a concept. For the concept “random error”, we collect multiple descriptions. The appropriate prompt is assigned based on the type of description provided.
  • Low Concept Bias:  ConceptPsy  covers all required concepts to provide a low-biased results for users.
  • Fine-grained Score: Instead of only providing an average score, we also provide chapter-level scores to reflect fine-grained strengths and weaknesses for developers.
  • An in-depth benchmark:Instead of aiming to cover as many subjects as possible, we focus on providing the comprehensive and high in-depth coverage within the field of psychology, encompassing all college-level subjects.
3.2 Data Collection
Subject selection.
[…]
Concepts collection.
We manually collect concepts based on the requirements of the 12 subjects outlined in National Post-graduate Entrance Examination. This task is undertaken by eight graduate students over the course of one month. Each concept is summarized from the tutorials of the exams. An example of the collected concepts is shown in Figure 5. For each concept, we summarize its description and manually determine which type of prompt (Calculation, Theory Understanding, Case Study) is suitable for generating questions about that concept (Prompts can be found in Appendix H). To enhance GPT-4’s understanding of the provided concepts, we supply not only the concepts themselves but also the name of the subject, and the primary and secondary headings they fall under.
Prompts design.
In the process of designing prompts to generate high-quality questions, we conduct a thorough analysis of question types found in the Professional Counselor Examination, categorizing them into three distinct groups: (1) calculation; (2) theory understanding; (3) case study (Appendix H). As depicted in Figure 5, while collecting concepts, we manually assign one or more prompt types to each concept. This assignment strategy ensures the relevance and diversity of the generated questions. To maintain a relatively even number of questions for each concept, we prompt GPT-4 to generate four questions per concept.
Figure 6: The review rules the question review process are as follows. Professional psychological annotators will filter, modify, and review the generated questions based on these requirements.
3.3 Questions Review
To ensure the quality of our dataset, we employ three professional psychological annotators to meticulously review each question. Our main objective is to verify the accuracy of the questions and to remove any that are found to be unreasonable. Each question is subjected to a thorough evaluation by at least one professional psychologist, who identifies any aspects that are not reasonable. When inaccuracies are discovered, all professional psychologists collaborate to discuss and agree upon a suitable and precise correction strategy. The criteria for reviewing questions are detailed in Figure 6. Throughout this process, the annotators check how relevant and accurate the knowledge points, the questions, and their connections are.
4 Experiments
4.1 Dataset Statistics
We carefully select 12 psychology subjects as shown in Table 5. For each concept, we generate 4 questions, leading to a total of 4573 high-quality questions after a thorough review and filtering process. Furthermore, as illustrated in Figure 2, we collect chapter names based on the syllabus of these subjects and assign chapter-level labels to each generated question to provide detailed insights about the strengths and weaknesses.
4.2 Setup
We evaluate a diverse set of strong models to ensure a comprehensive evaluation. More details about the models can be found in Appendix G. We conduct the evaluations using a 5-shot method. We set the temperature as 0, $top\_p$ as 1.0.
Clinical &
Counseling
Psy of
Personality
Abnormal
Psy
History of
Psy
General
Psy
Psy-
chometrics
Social
Psy
Management
Psy
Psychological
Statistics
Experimental
Psy
Developmental
Psy
Educational
Psy
Avg
GPT-4 0.83 0.86 0.88 0.81 0.84 0.80 0.89 0.87 0.79 0.85 0.82 0.86 0.84
GPT-3.5-Turbo 0.78 0.74 0.80 0.67 0.73 0.70 0.82 0.80 0.64 0.74 0.77 0.81 0.75
Chi-Alpaca2-7B 0.59 0.58 0.59 0.52 0.45 0.51 0.63 0.62 0.46 0.58 0.59 0.67 0.57
Llama-2-13B-Chat 0.66 0.62 0.65 0.56 0.54 0.57 0.66 0.63 0.48 0.57 0.58 0.66 0.60
Chatglm2-6B 0.71 0.63 0.68 0.63 0.62 0.62 0.69 0.64 0.54 0.64 0.66 0.72 0.65
Llama-2-70B-Chat 0.74 0.66 0.73 0.62 0.61 0.59 0.75 0.70 0.52 0.63 0.69 0.69 0.66
Mistral-7B-Ins 0.73 0.65 0.66 0.59 0.62 0.63 0.71 0.72 0.52 0.65 0.66 0.74 0.66
Baichuan2-7B-Chat 0.72 0.69 0.77 0.65 0.67 0.66 0.75 0.72 0.53 0.66 0.70 0.76 0.69
Baichuan2-13B-Chat 0.78 0.73 0.80 0.68 0.73 0.69 0.82 0.75 0.60 0.71 0.74 0.83 0.74
Mixtral-8x7B-Ins 0.82 0.76 0.78 0.65 0.73 0.70 0.83 0.78 0.71 0.73 0.74 0.79 0.75
Qwen1.5-7B-Chat 0.83 0.75 0.84 0.66 0.75 0.67 0.79 0.81 0.61 0.72 0.73 0.83 0.75
Internlm2-7B-Chat 0.86 0.75 0.84 0.68 0.78 0.7 0.82 0.82 0.62 0.75 0.78 0.81 0.77
Yi-6B-Chat 0.85 0.78 0.89 0.71 0.84 0.70 0.85 0.86 0.62 0.72 0.78 0.83 0.79
Qwen1.5-72B-Chat 0.87 0.74 0.89 0.80 0.81 0.77 0.88 0.87 0.38 0.77 0.83 0.85 0.79
Qwen1.5-14B-Chat 0.86 0.80 0.85 0.72 0.83 0.72 0.83 0.85 0.67 0.81 0.79 0.85 0.80
Yi-34B-Chat 0.92 0.84 0.91 0.81 0.86 0.78 0.88 0.90 0.69 0.84 0.86 0.86 0.85
Table 6: Performances on  ConceptPsy  with different LLMs.
Figure 7: Concept-level results for models more than 34B, in addition to GPT-3.5-Turbo.
4.3 Results
The main results of different LLMs on  ConceptPsy  are shown in Table 6. We also provide fine-grained results at Figure 7, 8 and 9.
Strong Chinese models achieve comparable overall results with GPT-4: In Table 6, surprisingly, some Chinese models achieve comparable performance close to that of GPT-4 on  ConceptPsy. Note that GPT-4 have the concepts as input during the question generation process but not in the evaluation stage. Overall, GPT-4 performed significantly better in subjects requiring more reasoning ability, such as Psychological Statistics (79% compared to 69%). However, in subjects requiring a deep understanding of the Chinese context and knowledge, such as Clinical & Counseling, Yi-34B-Chat performed better. This reflects the differences in abilities between the two models, indicating that GPT-4 might have insufficient training data in some areas of Chinese psychology, but its reasoning ability is much stronger.
Chinese models struggle with math and reasoning: Analysis from Figure 7, 8, and 9 reveals that Chinese models’ performance drops on concepts requiring math and reasoning, like “Mathematical Basis of Inferential Statistics” and “Non-Parametric Tests”. While these models excel in knowledge-based areas, they falter in math and reasoning segments. Mixtral-8x7B-Ins stands out by underperforming in knowledge questions but outshining others in math and reasoning. This suggests that despite high accuracy from training on extensive Chinese texts, Chinese models’ math and reasoning skills, especially in psychology, need enhancement.
Models show varied significant chapter performance: There’s notable variation in how models perform across chapters. Yi-34B-Chat, for example, despite leading in average accuracy, exhibits over 50% performance differences between chapters—excelling in basic concepts but lagging in math reasoning, even behind Mixtral-8x7B-Ins with its 10% lower average accuracy. This underscores  ConceptPsy’s role in offering detailed insights for model enhancement.
Variability in conceptual mastery across models from the same series: Although Qwen1.5-14B-Chat and Qwen1.5-72B-Chat achieve similar overall results, this does not mean their performance is similar across all chapters. As shown in Figures 7 and 8, Qwen1.5-72B-Chat performs much worse in chapters that focus on math and reasoning, like "Non-Parametric Tests" and "Parameter Estimation." While the overall accuracy suggests that both models perform similarly, the detailed results from  ConceptPsy  show that Qwen1.5-72B-Chat is generally better than Qwen1.5-14B-Chat in most chapters, except for a few math-related ones where it falls behind. This demonstrates how  ConceptPsy  can highlight specific strengths and weaknesses, helping developers make targeted improvements.
Subjects Pro Non-Pro
Clinical & Counseling Psychology 0.85 0.28
Psychology of Personality 0.82 0.24
Abnormal Psychology 0.89 0.32
History of Psychology 0.73 0.22
General Psychology 0.9 0.26
Psychometrics 0.75 0.3
Social Psychology 0.82 0.16
Management Psychology 0.8 0.36
Psychological Statistic 0.78 0.2
Experimental Psychology 0.82 0.32
Developmental Psychology 0.83 0.16
Educational Psychology 0.82 0.16
Avg 0.82 0.25
Table 7: Human evaluation results of  ConceptPsy. “Pro” and “Non-pro” represent Professional Counselor and Non-Professional Student respectively.
4.4 Ablation Study
Human baseline. To assess  ConceptPsy’s ability to distinguish between human expertise levels, we ask professional counselors and non-professional grad students to evaluate a subset of tasks. Table 7 shows significant performance differences (Appendix A).
The effect of linguist style. To investigate the effect of different linguistic styles of generated questions, we paraphrase each question using Qwen-70b-Chat evaluate LLMs on it. The results are shown in Table 8. There is a slight performance drop on paraphrased questions, potentially due to the lack of manual review for these paraphrased versions. Nonetheless, the consistency in model rankings and overall performance suggests that for multiple-choice questions, stylistic variations of the questions have limited effect.
5 Related Works
Chinese MMLU Benchmarks: While English language benchmarks continue to evolve (Hendrycks et al., 2020; Huang et al., 2023; Li et al., 2023; Zhong et al., 2023), Chinese MMLU benchmarks remain underdeveloped. The CLUE benchmark (Xu et al., 2020) is a widely used large-scale NLU benchmark for Chinese. AGIEval (Zhong et al., 2023) expands this with questions from various Chinese exams, and MMCU (Zeng, 2023) includes questions from diverse domains. The C-EVAL benchmark Huang et al. (2023) gathers questions from different educational levels and professional exams, using non-paper-based and simulation questions to prevent dataset leakage. ConceptMath (Wu et al., 2024) evaluates LLMs in a concept-wise manner but focuses solely on mathematics, is not a comprehensive MMLU benchmark, and was released more than three months after our work.
Benchmarks for LLMs in Psychology. Although psychology is a crucial domain for achieving artificial intelligence, there are currently few benchmarks in this field. CMMLU has collected hundreds of questions as "professional psychology" to evaluate a model’s knowledge understanding and reasoning capabilities. Other works (Yang et al., 2023b; Amin et al., 2023; Lamichhane, 2023) approach this from a mental health perspective, utilizing various classification tasks, including various emotions, suicidal tendencies, and more. PsyEval (Jin et al., 2023) further expands in the mental health field by adding tasks such as “Diagnosis Prediction”, but lacking comprehensive subjects in college-level psychology.
6 Discussion
Domain benchmarks are crucial for adapting LLMs. However, a comprehensive benchmark for assessing models’ knowledge and reasoning abilities in psychology is lacking. To address this gap, we introduce  ConceptPsy. In designing  ConceptPsy, we reveal concept bias issues in previous Chinese MMLU benchmarks and study concept coverage rates and performance variances. To tackle these issues, we propose a concept-wise question generation framework and provide chapter-level fine-grained results. Evaluation across various models demonstrates  ConceptPsy’s effectiveness in highlighting model strengths and weaknesses.
Limitations
We focused exclusively on concept bias within the Chinese MMLU dataset series, not exploring biases in other dimensions due to the labor-intensive nature of collecting the required concepts in a discipline, which necessitates trained annotators. Our experiments were thus confined to the field of psychology. Although our question generation method could be easily applied to other domains, due to copyright issues and the absence of corresponding multi-choice questions for many concepts, we opted to use GPT-4 for question generation instead of relying on human-designed questions.
Ethics Statement
We have thoroughly examined our data to ensure that there are no ethical issues. The data is generated by GPT-4, using concepts prescribed by the National Entrance Examination for Postgraduates. Furthermore, the generated data is subjected to rigorous scrutiny by professional psychologists to ensure its ethical soundness.
Appendix A Human Evaluation
We also provide human performance on the subset of  ConceptPsy. Specifically, we randomly sample 50 questions from each of 12 subjects, totaling 600 questions. A professional counselor independently completes all 600 questions, while ten non-experts are each assigned different subjects for evaluation. The results are shown in Table 7. The professional counselor significantly outperforms GPT-3.5, achieving an average score of 82%. However, the performance of non-expert undergraduates and graduate students is poor, averaging 25%, which is close to random. This underscores our benchmark’s demand for specialized knowledge in psychology and its effectiveness in distinguishing whether a model has mastered these professional concepts.
Appendix B The effect of Linguistic Style of generated questions
To investigate the effect of linguistic style, we paraphrase each question using Qwen-70b-Chat. We then evaluated the performance difference between its style and GPT-4’s style. The results are summarized in the table below:
Clinical &
Counseling
Psy of
Personality
Abnormal
Psy
History of
Psy
General
Psy
Psy-
chometrics
Social
Psy
Management
Psy
Psychological
Statistics
Experimental
Psy
Developmental
Psy
Educational
Psy
Avg
Mistral-7B-Instruct 0.71 0.63 0.64 0.58 0.57 0.59 0.69 0.71 0.53 0.64 0.62 0.69 0.63
gpt-3.5-Turbo 0.80 0.72 0.79 0.67 0.70 0.68 0.77 0.78 0.60 0.73 0.71 0.76 0.73
Yi-6B-Chat 0.80 0.75 0.87 0.67 0.80 0.70 0.83 0.83 0.58 0.69 0.76 0.86 0.76
Yi-34B-Chat 0.86 0.80 0.90 0.78 0.83 0.78 0.87 0.88 0.70 0.82 0.82 0.85 0.82
Table 8: Performances of models on questions paraphrased by Qwen-70b-Chat
In comparison to Table 6, models perform slightly worse on paraphrased questions generated by Qwen-70b-Chat. This may be because the paraphrased questions are not manually reviewed. However, despite the style changes, the model rankings and overall performance remained consistent. This indicates that for multiple-choice questions, the style of the questions has minimal impact on evaluation.
Subjects Win Rate Tie Rate
Clinical & Counseling Psychology 0 0.8
Psychology of Personality 0 0.9
Abnormal Psychology 0 0.8
History of Psychology 0.1 0.8
General Psychology 0.1 0.9
Psychometrics 0.1 0.7
Social Psychology 0.82 0.16
Management Psychology 0.1 0.8
Psychological Statistic 0.4 0.5
Experimental Psychology 0.2 0.7
Developmental Psychology 0.4 0.4
Educational Psychology 0.3 0.7
Avg 0.27 0.62
Table 9: We hire two professional counselor to compete the quality of generated questions and human-designed questions.
Appendix C Competition with Human-designed Questions
We randomly selected 15-20 questions for each subject and an equivalent number of questions from real exams available online. We shuffle the order of the questions, presenting one generated and one real question side by side, and ask two psychological annotators to judge which was better or if there was a tie. The averaged results are shown in Table 9. The “Win Rate” indicates the proportion of instances where our generated questions were rated better than the real ones. As can be seen, the majority of the generated questions matched or even surpassed the performance of actual exam questions. This success can be attributed to two factors: 1) our prompts are a carefully designed summary of the question types found in Chinese psychological professional examinations; 2) GPT-4’s robust capabilities in synthetic data.
Appendix D Chapter-level Statistic of ConceptPsy
In Table 10 and Table 11, we present the chapters of each subject along with the corresponding number of concepts and questions.
Subject Chapter # C # Q
Clinical & Counseling Psychology History of Clinical and Counseling Psychology 4 13
Basic concepts of psychotherapy and counseling 3 7
Characteristics of the therapeutic relationship and its influences 10 32
Work Ethics in Clinical and Counseling Psychology 5 16
Research Methods in Clinical and Counseling Psychology 8 25
Theory and Practice of Counseling and Psychotherapy 25 63
Psychology of Personality Basic Concepts of Personality Psychology 12 43
Psychoanalytic School 23 84
Behaviorist school of personality theory 10 37
Cognitive School 9 33
Humanistic School 11 40
Personality Trait Theory 18 57
Biological School 4 13
Positive Psychology 4 11
Abnormal Psychology Basic Concepts of Abnormal Psychology 17 60
Anxiety Disorders 26 81
Mood Disorders 7 19
Eating Disorders 6 18
Personality Disorders 14 40
Substance Dependence 10 28
Childhood Mental Disorders 8 22
History of Psychology The Origins and Establishment of Western Psychology 17 62
Psychology of Consciousness 40 154
Behaviourist Psychology 21 82
Psychoanalysis 25 93
Cognitive Psychology 13 48
Humanistic Psychology 8 30
History of Chinese Psychology 1 3
General Psychology Basic Concepts in General Psychology 15 48
Biological Basis of Mind and Behaviour 9 33
Consciousness and Attention 15 51
Sensation 15 46
Perception 17 60
Memory 16 53
Thinking 21 67
Speech 7 21
Moods and Emotions 15 50
Motivation, Needs & Will 15 52
Competencies 15 50
Personality Theory 22 74
Psychometrics Basic Concepts in Measurement Psychology 12 41
Classical Measurement Theory 24 91
Basic Concepts of Psychological Testing 40 105
Commonly Used Psychological Tests 39 129
Table 10: Details of the number of Concepts and Questions in each Chapter.
Subject Chapter # C # Q
Social Psychology History of Social Psychology 40 134
Social Thinking 48 163
Social Relations 27 94
Social Influence 24 79
Basic Concepts of Social Psychology 30 89
Management Psychology Management Philosophy 18 65
Organizational Motivation 18 80
Leadership Theory 18 65
Organizational Theory 33 105
Psychological Statistics Basic Concepts in Statistical Psychology 3 12
Statistical Charts 5 15
Concentration 3 13
Measures of Variation 7 34
Correlation 2 5
Mathematical Basis of Inferential Statistics 16 65
Parameter Estimation 13 28
Hypothesis Testing 13 53
Applied Computing in Statistical Psychology 21 62
Chi-Square Tests 2 5
Non-Parametric Tests 2 8
Preliminary Multivariate Statistical Analysis 2 11
Experimental Psychology Basic Concepts of Experimental Psychology 3 12
Variables in Psychological Experiments 23 75
Design of Psychological Experiments 12 38
Reaction Time Method 9 33
Psychophysical Methods 40 97
Major Psychological Experiments 47 158
Developmental Psychology Basic Concepts in Developmental Psychology 5 24
Basic Theories of Psychological Development 19 90
Biological Basis of Psychological Development and Fetal Development 9 33
Intelligence 9 34
Emotions 8 34
Early Childhood Psychological Development 36 114
Child Psychological Development 27 90
Adolescent Psychological Development 18 60
Psychological Development in Adulthood 27 101
Educational Psychology Basic Concepts of Educational Psychology 6 18
General Psychology of Learning 6 21
Major Theories of Learning 29 100
Categorical Learning Psychology 25 69
Table 11: Details of the number of Concepts and Questions in each Chapter (Cont.)
Appendix E Experiments Details of Calculating Concept Coverage Rate
To calculate the concept coverage rate, the challenge lies in obtaining the required concepts in a subject. For advanced math, we first collect its chapters and the concepts under different chapters based on relevant exam requirements. We initially prompt GPT-4 to classify each question into different chapters, then prompt GPT-4 to further classify the question into specific instance-level concepts. We calculate the concept coverage rate and performance variance based on the concepts covered by these questions. This hierarchical classification allows for more accurate categorization.
For psychology in CMMLU, due to the vast number of concepts, prompting GPT-4 with these concepts in the input prompt each time is too expensive, so we use chapter-level concepts. We collect 84 chapter names according to the requirements, and classify each question into one or more chapters.
For other subjects, manually collecting concepts is too costly, and it’s impossible to collect concepts for each subject. We prompt GPT-4 to generate a 3-level syllabus for the subject as the required concepts. We then prompt GPT-4 to categorize each question into one or more first-level headings in the syllabi. We further classify them into the second- and third-level headings under the selected first-level headings.
Appendix F More Results on Fine-grained Performance
We provide more chapter-level results in this Figure 8, 9.
Figure 8: Concept-level results for models wi 14 Billion parameters
Figure 9: Concept-level results for some other models
Appendix G Details About Evaluated Models
Model Name Model Code/API
Chinese-Alpaca-2-7BCui et al. (2023) hfl/chinese-alpaca-2-7b
Chatglm2-6BDu et al. (2022) THUDM/chatglm2-6b
Llama-2-13B-ChatTouvron et al. (2023) meta-llama/Llama-2-13b-chat-hf
Llama-2-70B-ChatTouvron et al. (2023) meta-llama/Llama-2-70b-chat-hf
Baichuan2-7B-ChatYang et al. (2023a) baichuan-inc/Baichuan2-7B-Chat
Baichuan2-13B-ChatYang et al. (2023a) baichuan-inc/Baichuan2-13B-Chat
Mistral-7B-Instruct-v0.2Jiang et al. (2023) mistralai/Mistral-7B-Instruct-v0.2
Mixtral-8x7B-Instruct-v0.1Jiang et al. (2024) mistralai/Mixtral-8x7B-Instruct-v0.1
Qwen1.5-7B-ChatBai et al. (2023b) Qwen/Qwen1.5-7B-Chat
Qwen1.5-72B-ChatBai et al. (2023b) Qwen/Qwen1.5-72B-Chat
Qwen1.5-14B-ChatBai et al. (2023b) Qwen/Qwen1.5-14B-Chat
Internlm2-7B-ChatCai et al. (2024) internlm/internlm2-chat-7b
Yi-6B-ChatAI et al. (2024) 01-ai/Yi-6B-Chat
Yi-34B-ChatAI et al. (2024) 01-ai/Yi-34B-Chat
GPT-3.5-TurboOpenAI (2022) Azure api: gpt-35-turbo
Table 12: Model code/API of our evaluated models.
The evaluated models include Chinese-Alpaca-2-7BCui et al. (2023), Chatglm-6BDu et al. (2022), Chatglm2-6BDu et al. (2022), Llama-2-13B-ChatTouvron et al. (2023), Llama-2-70B-ChatTouvron et al. (2023), Baichuan2-7B-ChatYang et al. (2023a), Baichuan2-13B-ChatYang et al. (2023a), Mistral-7B-Instruct-v0.2Jiang et al. (2023), Mixtral-8x7B-Instruct-v0.1Jiang et al. (2024), Qwen1.5-7B-ChatBai et al. (2023b), Qwen1.5-72B-ChatBai et al. (2023b), Qwen1.5-14B-ChatBai et al. (2023b), Internlm2-7B-ChatCai et al. (2024), Yi-6B-ChatAI et al. (2024), and Yi-34B-ChatAI et al. (2024). The model codes can be found in Table 12 . We evaluate these models with vLLM (Kwon et al., 2023).
Appendix H Prompts for Questions Generation
We totally design four distinct prompts to steer GPT-4 in generating questions based on provided knowledge points. In designing these prompts, we have taken into account several guidelines. Firstly, the generated questions should exhibit a high level of difficulty and complexity. Secondly, while generating questions, GPT-4 should primarily rely on the given knowledge points, but it can also incorporate its inherent psychological knowledge. Thirdly, each knowledge point unit may yield multiple questions, but the content, type, or perspective of the questions should be distinct.
Figures 10, 11, 12, and 13 show the specific prompts that we have inputted into GPT-4. These prompts have been meticulously designed, with each one being tailored to control the generation of a different kind of question. The first three prompts correspond to generating Theory understanding, Case study, and Calculation type questions, respectively, while the last one encompasses all of the aforementioned types. We choose the appropriate prompt for each knowledge point based on its content.
Figure 10: Theory understanding
Figure 11: Case study
Figure 12: Calculation
Figure 13: Multiple types prompt
Figure 14: The question generation prompt template (translated in English), which is primarily designed for generating the type of calculation questions.
Figure 15: An example of prompts in few-shot setting. The black text is what we feed into model, while the red text is the response completed by model. The English translation for the Chinese input is provided in the purple text, which is not included in the actual prompt.